OpenStudy (anonymous):

show all work. Find the product: (2/3x^2-3/5)(2/3x -1/2)

OpenStudy (anonymous):

\[( 2/3x^2 - 3/5 ) ( 2/3x - 1/2 ) \] \[2/3x^2 * 2/3x = 4/9x^3\] \[2/3x^2 * - 1/2 = -2/6 x^2\] \[-3/5*2/3x = -6/15 x = - 3/5x\] \[-3/5*-1/2 = 3/10\] answer: \[4/9x^3 - 1/3 x^2 - 3/5x + 3/10\]